Present and Future Values of Single Cash Flows for Different Interest Rates Use both the TVM equations and a financial calculator to find the following values. Round your answers to the nearest cent. (Hint: Using a financial calculator, you can enter the known values and then press the appropriate key to find the unknown variable. Then, without clearing the TVM register, you can "override" the variable that changes by simply entering a new value for it and then pressing the key for the unknown variable to obtain the second answer. This procedure can be used in parts b and d, and in many other situations, to see how changes in input variables affect the output variable.) An initial $700 compounded for 10 years at 3.8 percent. An initial $700 compounded for 10 years at 7.6 percent. The present value of $700 due in 10 years at a 3.8 percent discount rate. The present value of $700 due in 10 years at a 7.6 percent discount rate.
Answer #1

1). An initial $700 compounded for 10 years at 3.8%

PV =700; N = 10; rate = 3.8%; PMT = 0, CPT FV. FV = 1,016.42

FV = PV*(1+rate)^N = 700*(1+3.8%)^10 = 1,016.42

2). An initial $700 compounded for 10 years at 7.6%

PV =700; N = 10; rate = 7.6%; PMT = 0, CPT FV. FV = 1,456.20

FV = PV*(1+rate)^N = 700*(1+7.6%)^10 = 1,456.20

3). The present value (PV) of $700 due in 10 years at a 3.8%

FV = 700; N = 9 (since it is annuity due, it will be paid at the beginning of year 10 or end of year 9); rate = 3.8%; PMT = 0, CPT PV. PV = 500.41

PV of annuity due = FV/(1+ rate)^9 = 700/(1+3.8%)^9 = 500.41

4). The present value (PV) of $700 due in 10 years at a 7.6%

FV = 700; N = 9; rate = 7.6%; PMT = 0, CPT PV. PV = 362.07

PV of annuity due = FV/(1+rate)^9 = 700/(1+7.6%)^9 = 362.07